# High-performance Effective Scientific Error-bounded Lossy Compression with Auto-tuned Multi-component Interpolation

### v2

@@ Abstract.
Error-bounded lossy compression has been identified as a promising solution for significantly reducing scientific data volumes upon users’ requirements on data distortion. For the existing scientific error-bounded lossy compressors, some of them (such as SPERR and FAZ) can reach fairly high compression ratios and some others (such as SZx, SZ, and ZFP) feature high compression speeds, but they rarely exhibit both high ratio and high speed meanwhile. In this paper, we propose HPEZ with newly-designed interpolations and quality-metric-driven auto-tuning, which features significantly improved compression quality upon the existing high-performance compressors, meanwhile being exceedingly faster than high-ratio compressors. The key contributions lie in the following points: (1) We develop a series of advanced techniques such as interpolation re-ordering, multi-dimensional interpolation, and natural cubic splines to significantly improve compression qualities with interpolation-based data prediction. (2) The auto-tuning module in HPEZ has been carefully designed with novel strategies, including but not limited to block-wise interpolation tuning, dynamic dimension freezing, and Lorenzo tuning. (3) We thoroughly evaluate HPEZ compared with many other compressors on six real-world scientific datasets. Experiments show that HPEZ outperforms other high-performance error-bounded lossy compressors in compression ratio by up to 140% under the same error bound, and by up to 360% under the same PSNR. In parallel data transfer experiments on the distributed database, HPEZ achieves a significant performance gain with up to 40% time cost reduction over the second-best compressor.
copyright: none
@@ 1. Introduction
The gigantic scale and exceptionally intense computation power of modern supercomputers have empowered the exascale scientific simulation applications to generate tremendous amounts of data in short periods, bringing up significant burdens for distributed scientific databases and cloud data centers. For instance, A one-trillion particle Hardware/Hybrid Accelerated Cosmology Code (HACC) (Habib et al., 2016) can harness approximately 22PB output data in a single simulation (Viswanash et al., 2019), and Community Earth System Model (CESM) (Kay and et al., 2015) simulation may generate 2.5PB data for a simulation task (Straatsma et al., 2017). To this end, error-bounded lossy compression techniques have been developed for those scientific data, and they have been recognized as the most proper strategy to manage the extremely large amount of data. The advantage of error-bounded lossy compression is primarily two-fold. On the one hand, it can reduce the original data to an incredibly shrunken size which is much smaller than the compressed data size generated by a lossless compressor. On the other hand, the error-bounded lossy compression can constrain the point-wise data distortion strictly upon the users’ requirements. Existing state-of-the-art error-bounded lossy compressors in diverse archetypes, such as prediction-based model – SZ3 (Zhao et al., 2021; Liang et al., 2022) and QoZ (Liu et al., 2022), transform-based model – ZFP (Lindstrom, 2014) and SPERR (Li et al., 2023), and dimension-reduction-based model – TTHRESH (Ballester-Ripoll et al., 2019), have been widely adopted in many use cases in practice.
@@
For modern scientific databases and cloud data centers which often involve multiple sites over a wide area network (WAN), the extremely large amount of raw data costs an unacceptable time to transfer between machines. Therefore, data compressors are critical for efficient data transfer because transferring compressed data will significantly reduce the time cost, as confirmed by prior research (Li et al., 2021; Liu et al., 2023a). In this case, compression ratios and speeds are both critical for achieving high data transfer throughput.
However, designing a versatile error-bounded lossy compressor that delivers high compression ratios with sufficient performance (specifically, the ’performance’ means speed) is quite challenging. On one hand, to reach a high compression performance, general techniques have to perform relatively simple data transform (Lindstrom, 2014) or prediction within short-range areas (Liang et al., 2018; Zhao et al., 2021; Ainsworth et al., 2018), which cannot take advantage of long-range data correlations, thus leading to very limited compression ratios inevitably. On the other hand, to reach a high compression ratio, general techniques are applying sophisticated techniques such as wavelet transform on the full data input (Liu et al., 2023b; Li et al., 2023) or higher-order SVD (Ballester-Ripoll et al., 2019), which suffer from very expensive operations inevitably, conflicting with our high-performance objective. As such, we must design more compact and effective data operation methods with relatively low computational costs, featuring high speed meanwhile yielding comparable compression ratios compared to the existing high-ratio compression techniques.
In order to design an error-bounded compressor that features both high compression ratios and satisfactory speeds, we propose an optimized quality-metric-driven error-bounded lossy compressor (called HPEZ) by developing a brand-new auto-tuning strategy and an anchor-based level-wise hybrid interpolation predictor. Integrating extensively optimized interpolation predictors and auto-tuning modules, HPEZ attains far better compression ratios and lower distortions than other high-performance error-bounded lossy compressors with limited compression speed degradation. HPEZ substantially outperforms high-ratio compressors in terms of speed. It achieves optimized throughput performance in a variety of use cases such as parallel data transfer for large (distributed) databases. We attribute our contributions as follows:
@@
  • We develop a series of optimization strategies including block-wise interpolation tuning, dynamic dimension freezing, and Lorenzo tuning, which can substantially improve the adaptability of the auto-tuning module for optimizing the compression across a broad spectrum of inputs.
@@ 2. Related Work
The prediction-based compressors use data prediction techniques, like linear regression (Liang et al., 2018) and dynamic spline interpolations (Zhao et al., 2021), to anticipate the data points. Well-known examples are SZ2 (Liang et al., 2018) and SZ3 (Liang et al., 2022; Zhao et al., 2021). Transform-based compressors, on the other hand, use data transformations to de-correlate the data, then switch to compress the generated coefficients which are more compressible. ZFP (Lindstrom, 2014), for example, is a typical example that employs exponent alignment, orthogonal discrete transform, and embedded encoding. SPERR (Li et al., 2023), a more recent work, leverages wavelet transform for data compression. Dimension-reduction-based compressors apply dimension reduction techniques, with (high-order) singular vector decomposition (SVD) being a case in point (for instance, TTHRESH (Ballester-Ripoll et al., 2019)). Neural-network-based compressors (Liu et al., 2021a; Glaws et al., 2020; Liu et al., 2021b; Hayne et al., 2021) utilize neural network models like the autoencoder family (Bank et al., 2020; Kingma and Welling, 2013; Kolouri et al., 2018), however, the speeds of them and relatively quite slow.
@@ 3.2. Determining the Best-fit Compressor Archetype for HPEZ
As mentioned before, our proposed compressor should exhibit both good rate-distortion and relatively high speeds. To this end, we need to investigate existing scientific error-bounded lossy compressors to identify the best-fit compressor archetype for our design. The categorization of compressors is priorly discussed in Section 1 and Section 2, but to conduct a deeper analysis here we split the existing state-of-the-art error-bounded compressors into more types according to their design details:
@@
  • Interpolation-based: Leveraging hierarchical interpolations for prediction-based data compression (Zhao et al., 2021; Liu et al., 2022).
@@ 4. HPEZ Design Overview
  • Step 4: Huffman encoding: The quantized prediction errors are further encoded with Huffman encoding. A more concentrated distribution of quantization errors will lower the encoded tree size, therefore the reduction of prediction error is key to improving the compression ratio.
@@ 5. HPEZ Interpolation-based Predictor
In this section, we describe the details of our fine-tuned multi-component interpolation-based data predictor for HPEZ. Compared to the existing interpolation-based predictors, the HPEZ interpolation-based predictor projects a significant improvement over them, especially attributed to several new components we designed and proposed. These components can obtain a significantly improved prediction accuracy, thus leading to much better rate distortions in the compression. Those new designs together with the existing interpolation designs will be described in the rest of this section and will get auto-tuned for optimization of compression quality (to be detailed in Section 6).
@@ 5.1. Overview of Interpolation-based Prediction
  • The interpolations are done level by level, from large strides (half of the anchor stride) to small strides (1).
@@ 5.2. Spline Interpolation Formulas
The three spline functions $f_{1}$, $f_{2}$, and $f_{3}$ have scopes of [$i-$3,$i-$1], [$i-$1,$i+$1], and [$i+$1,$i+$3], respectively. The zero-order, first-order, and second-order interpolation conditions are shown as follows:
@@
(7) $\vspace{-2mm}\begin{array}[]{l}f_{1}^{{}^{\prime\prime}}(i-3)=0;f_{3}^{{}^{\prime\prime}}(i+3)=0\\ \end{array}$
@@
The existing 1D-style interpolation actually has not fully exploited the multi-dimensional continuity and smoothness of input data arrays, because all the interpolations are constricted in a single dimensional direction. To address this limitation, we propose a new interpolation paradigm for HPEZ called multi-dimensional spline interpolation, which can take better advantage of data correlation across multiple dimensions. As shown in Figure 4 (b), the multi-dimensional spline interpolation initially performs the 1D interpolations for some data points as there are only 1D neighbors at the moment, then it performs 2D interpolations for the remaining data points that already have neighbors in two dimensions. The multi-dimensional spline interpolation is symmetric across all the dimensions, meaning that it does not need a selection of dimensional order.
@@
We feature the HPEZ multi-dimensional interpolation as follows. For each data point $x$, suppose $X_{i}\ (1\leq i\leq n)$ are all the available 1D interpolation results for predicting $x$ (which can either be linear interpolation or cubic interpolation), the multi-dimensional interpolation result $X^{{}^{\prime}}$ is a linear-combination of $X_{i}$:
@@
As such, we have proved that, if the $\{\alpha_{i}\}$ is selected based on Eq. 12, the prediction error variance of the multi-dimensional interpolation $X^{{}^{\prime}}$ will be no larger than each of the 1D-style interpolation $X_{i}$ according to Eq. 11. So, the average L-1 prediction error will also be minimized (which can be easily deducted). ∎
@@ 5.4. Interpolation Re-ordering
After the proposal of natural cubic spline interpolation and multi-dimensional spline interpolation, HPEZ also introduces interpolation re-ordering, which improves both prediction accuracy and prediction speed. It includes two aspects: the fast-varying-first interpolation and same-level cubic interpolation.
@@ 5.4.2. Same-level cubic interpolation
We develop a new same-level cubic interpolation in HPEZ, which can further improve prediction accuracy. In the traditional interpolation design (Liu et al., 2022; Zhao et al., 2021), at each interpolation level, the neighbor points of each data point to be interpolated are limited on the higher levels (interpolation levels with larger strides). For the 1D cubic spline interpolation applied on a data point with stride $s$, 4 neighbor points with distance $s$ and $3s$ are used, which have been predicted on the higher interpolation levels. As shown in Figure 6 (note that $s$ is the distance between each closest white and blue point), the first row shows this interpolation method, in which all the white data points (on the current interpolation level) are predicted by the blue data points (on higher interpolation levels). If we are able to include more neighbors for each point (for example, the 2 white points with a distance of $2s$ to it), the prediction accuracy can be improved. As illustrated in the 2nd and 3rd rows of Figure 6, instead of traversing through all the white data points in one step, HPEZ splits the 1D cubic spline interpolation into 2 steps. In the first step (the second row of Figure 6), half of the white points are interpolated by inter-level interpolation (the existing interpolation) with 4 neighbor points. In the second round, the rest half of the white points are interpolated by the same-level interpolation with 6 neighbor points for each, including blue points (on higher interpolation levels) and green points (on the current interpolation level). With this new interpolation, half of the data points are predicted with two more neighbor points to achieve better prediction accuracy. Similar to the deductions in Section 5.2, for a data point $p_{i}$, with its 6 neighbor points $d_{i-3}$, $d_{i-2}$, $d_{i-1}$, $d_{i+1}$, $d_{i+2}$, and $d_{i+3}$ the same-level cubic spline interpolation formula would be the following two. Eq. 13 is for the not-a-knot cubic spline and Eq. 14 is for the natural cubic spline. The same strategy can also be extended to the multi-dimensional interpolation, splitting it into 2 steps each with halved data points.
@@ 6.2. Global Interpolation Tuning
  • New in HPEZ: Whether applying inner-level interpolation or not (only for cubic interpolation);
Similar to QoZ, the sampled data from the data sampling process are used for performing compression tests with all the available interpolation configurations. Then, HPEZ selects the interpolation configuration with the lowest average absolute prediction error as the final tuning result.
@@ 6.3. Dynamic Dimension Freezing
The dynamic dimension freezing in HPEZ is designed to avoid inaccurate interpolation predictions along non-smooth dimensions. For a multi-dimensional input data array, it may present fine smoothness along some of its dimensions but present bad smoothness along the other dimensions. In those cases, both the 1D-style and multi-dimensional interpolation will fail in achieving high prediction accuracy as they will involve interpolations along non-smooth directions. The dimension freezing is that, given one dimension, HPEZ sets anchor points along those dimensions with stride 1 (without intervals) and never performs interpolations along those dimensions. Figure 8 uses the interpolation on a 3D data block as an example of dimension freezing. For a clear view, only the 1D interpolations are shown. Figure 8 (a) is the normal 1D interpolations without a frozen dimension, and Figure 8 (b) is the 1D interpolations with a dimension frozen, in which no interpolations are made along the frozen dimension. With this dynamic strategy, HPEZ is free of relying on the smoothness of data along all dimensions to optimize its compression ratio. According to our experimental results, compared to the highly improved prediction accuracy and greatly reduced quantization bin size, the storage overhead for additional anchor points is trivial. To determine whether to freeze a dimension and which dimension should be frozen, the auto-tuning module of HPEZ first specifies the most non-smooth dimension in the input data array in the statistical analysis (Section 6.1), then separately tunes 2 optimized interpolation configurations with/without this dimension frozen. If freezing this dimension presents a better compression ratio, HPEZ will freeze this dimension.

### v3

@@ Abstract.
Error-bounded lossy compression has been identified as a promising solution for significantly reducing scientific data volumes upon users’ requirements on data distortion. For the existing scientific error-bounded lossy compressors, some of them (such as SPERR and FAZ) can reach fairly high compression ratios and some others (such as SZx, SZ, and ZFP) feature high compression speeds, but they rarely exhibit both high ratio and high speed meanwhile. In this paper, we propose HPEZ (a.k.a. QoZ 2.0) with newly designed interpolations and quality-metric-driven auto-tuning, which features significantly improved compression quality upon the existing high-performance compressors, meanwhile being exceedingly faster than high-ratio compressors. The key contributions lie as follows: (1) We develop a series of advanced techniques such as interpolation re-ordering, multi-dimensional interpolation, and natural cubic splines to significantly improve compression qualities with interpolation-based data prediction. (2) The auto-tuning module in HPEZ has been carefully designed with novel strategies, including but not limited to block-wise interpolation tuning, dynamic dimension freezing, and Lorenzo tuning. (3) We thoroughly evaluate HPEZ compared with many other compressors on six real-world scientific datasets. Experiments show that HPEZ outperforms other high-performance error-bounded lossy compressors in compression ratio by up to 140% under the same error bound, and by up to 360% under the same PSNR. In parallel data transfer experiments on the distributed database, HPEZ achieves a significant performance gain with up to 40% time cost reduction over the second-best compressor.
error-bounded lossy compression, interpolation, scientific database
copyright: licensedusgovmixedjournal: PACMMODjournalyear: 2024journalvolume: 2journalnumber: 1 (SIGMOD)article: 4publicationmonth: 2doi: 10.1145/3639259ccs: Information systems Data compressionccs: Theory of computation Data compressionccs: Mathematics of computing Interpolation
@@ 1. Introduction
The gigantic scale and exceptionally intense computation power of modern supercomputers have empowered the exascale scientific simulation applications to generate tremendous amounts of data in short periods, bringing up significant burdens for distributed scientific databases and cloud data centers. For instance, A one-trillion particle Hardware/Hybrid Accelerated Cosmology Code (HACC) (Habib et al., 2016) can harness approximately 22PB output data in a single simulation, and Community Earth System Model (CESM) (Kay and et al., 2015) simulation may generate 2.5PB data for a simulation task (Straatsma et al., 2017). To this end, error-bounded lossy compression techniques have been developed for those scientific data, and they have been recognized as the most proper strategy to manage the extremely large amount of data. The advantage of error-bounded lossy compression is primarily two-fold. On the one hand, it can reduce the original data to an incredibly shrunken size which is much smaller than the compressed data size generated by a lossless compressor. On the other hand, the error-bounded lossy compression can constrain the point-wise data distortion strictly upon the users’ requirements. Existing state-of-the-art error-bounded lossy compressors in diverse archetypes, such as prediction-based model – SZ3 (Zhao et al., 2021; Liang et al., 2022) and QoZ (Liu et al., 2022), transform-based model – ZFP (Lindstrom, 2014) and SPERR (Li et al., 2023), and dimension-reduction-based model – TTHRESH (Ballester-Ripoll et al., 2019), have been widely adopted in many use cases in practice.
@@
For modern scientific databases and cloud data centers which often involve multiple sites over a wide area network (WAN), the extremely large amount of raw data costs an unacceptable time to transfer between machines. Therefore, data compressors are critical for efficient data transfer because transferring compressed data will significantly reduce the time cost, as confirmed by prior research (Li et al., 2021; Liu et al., 2023a). In this case, compression ratios and speeds are both critical for achieving high data transfer throughput. However, designing a versatile error-bounded lossy compressor that delivers high compression ratios with sufficient performance (i.e. speed) is quite challenging. On one hand, to reach a high compression performance, general techniques have to perform relatively simple data transform (Lindstrom, 2014) or prediction within short-range areas (Liang et al., 2018; Zhao et al., 2021; Ainsworth et al., 2018), which cannot take advantage of long-range data correlations, thus leading to very limited compression ratios inevitably. On the other hand, to reach a high compression ratio, general techniques are applying sophisticated techniques such as wavelet transform on the full data input (Liu et al., 2023b; Li et al., 2023) or higher-order SVD (Ballester-Ripoll et al., 2019), which suffer from very expensive operations inevitably, conflicting with our high-performance objective. As such, we must design more compact and effective data operation methods with relatively low computational costs, featuring high speed, meanwhile yielding comparable compression ratios compared to the existing high-ratio compression techniques.
In order to design an error-bounded compressor that features both high compression ratios and satisfactory speeds, we propose an optimized quality-metric-driven error-bounded lossy compressor HPEZ (also known as QoZ 2.0 as it serves as the second major version of QoZ) by developing a brand-new auto-tuning strategy and an anchor-based level-wise hybrid interpolation predictor. Integrating extensively optimized interpolation predictors and auto-tuning modules, HPEZ attains far better compression ratios and lower distortions than other high-performance error-bounded lossy compressors with limited compression speed degradation. HPEZ substantially outperforms high-ratio compressors in terms of speed. It achieves optimized throughput performance in a variety of use cases such as parallel data transfer for large (distributed) databases. We attribute our contributions as follows:
@@
  • We develop a series of optimization strategies including block-wise interpolation tuning, dynamic dimension freezing, and Lorenzo tuning, which can substantially improve the adaptability of the auto-tuning for the compression across a broad spectrum of inputs.
@@ 2. Related Work
The prediction-based compressors use data prediction techniques, like linear regression (Liang et al., 2018) and dynamic spline interpolations (Zhao et al., 2021), to anticipate the data points. Well-known examples are SZ2 (Liang et al., 2018) and SZ3 (Liang et al., 2022; Zhao et al., 2021). Transform-based compressors, on the other hand, use data transformations to de-correlate the data, then switch to compress the more compressible transformed coefficients. ZFP (Lindstrom, 2014), for example, is a typical example that employs exponent alignment, orthogonal discrete transform, and embedded encoding. SPERR (Li et al., 2023), a more recent work, leverages wavelet transform for data compression. Dimension-reduction-based compressors apply dimension reduction techniques, with (high-order) singular vector decomposition (SVD) being a case in point (for instance, TTHRESH (Ballester-Ripoll et al., 2019)). Neural-network-based compressors (Liu et al., 2021a; Glaws et al., 2020; Liu et al., 2021b; Hayne et al., 2021) utilize neural network models like the autoencoder family (Bank et al., 2020; Kingma and Welling, 2013; Kolouri et al., 2018), however, the speeds of them and relatively quite slow.
@@ 3.2. Determining the Best-fit Compressor Archetype for HPEZ
As mentioned before, our proposed compressor should exhibit both good rate-distortion and relatively high speeds. To this end, we need to investigate existing scientific error-bounded lossy compressors to identify the best-fit compressor archetype for our design. The categorization of compressors is priorly discussed in Section 1 and Section 2, but to conduct a deeper analysis here we categorize the existing compressors into more types according to their designs:
@@
  • Interpolation-based: Leveraging interpolations for prediction-based data compression (Zhao et al., 2021; Liu et al., 2022).
@@ 4. HPEZ Design Overview
  • Step 4: Huffman encoding: The quantized prediction errors acquired from Step 3 are further encoded with Huffman encoding. A more concentrated distribution of quantization errors will lower the encoded tree size, therefore the reduction of prediction error is key to improving the compression ratio.
@@ 5. HPEZ Interpolation-based Predictor
In this section, we describe the details of our fine-tuned multi-component interpolation-based data predictor for HPEZ. Compared to the existing interpolation-based predictors, the HPEZ interpolation-based predictor projects a significant improvement over them, attributed to several new components we designed and proposed. These components can obtain a significantly improved prediction accuracy, thus leading to much better rate distortions in the compression. Those new designs together with the existing interpolation designs will be described in the rest of this section and will get auto-tuned for optimization of compression quality (to be detailed in Section 6).
@@ 5.1. Overview of Interpolation-based Prediction
  • The interpolations are done hierarchically (level by level), from large strides (half of the anchor stride) to small strides (1).
@@ 5.2. Spline Interpolation Formulas
The spline functions $f_{1}$, $f_{2}$, and $f_{3}$ have scopes of [$i-$3,$i-$1], [$i-$1,$i+$1], and [$i+$1,$i+$3], respectively. The zero-order, first-order, and second-order interpolation conditions are shown as follows:
@@
(7) $\vspace{-2mm}\begin{array}[]{l}f_{1}^{{}^{\prime\prime}}(i-3)=0;f_{3}^{{}^{\prime\prime}}(i+3)=0\\ \end{array}\vspace{1mm}$
@@
Actually, The existing 1D-style interpolation has not fully exploited the multi-dimensional continuity and smoothness of input data arrays, because all the interpolations are constricted in a single-dimensional direction. To address this limitation, we propose a new interpolation paradigm for HPEZ called multi-dimensional spline interpolation, which can take better advantage of data correlation across multiple dimensions. As shown in Figure 4 (b), the multi-dimensional spline interpolation initially performs the 1D interpolations for some data points as there are only 1D neighbors at the moment, then it performs 2D interpolations for the remaining data points that already have neighbors in two dimensions. The multi-dimensional spline interpolation is symmetric across all the dimensions, meaning that it does not need a selection of dimensional order.
@@
We feature the HPEZ multi-dimensional interpolation as follows. For each data point $x$, suppose $X_{i}\ (1\leq i\leq n)$ are all the available 1D interpolation results for predicting $x$ (which can either be linear interpolation or cubic interpolation and are along all dimensions), the multi-dimensional interpolation result $X^{{}^{\prime}}$ is a linear-combination of $X_{i}$:
@@
As such, we have proved that, if the $\{\alpha_{i}\}$ is selected based on Eq. 12, the prediction error variance of the multi-dimensional interpolation $X^{{}^{\prime}}$ will be no larger than each of the 1D-style interpolation $X_{i}$ according to Eq. 11. So, the average L-1 prediction error will also be minimized. ∎
@@ 5.4. Interpolation Re-ordering
After the proposal of natural cubic spline and multi-dimensional interpolation, HPEZ also introduces interpolation re-ordering, which improves both prediction accuracy and prediction speed. It includes two aspects: the fast-varying-first interpolation and same-level cubic interpolation.
@@ 5.4.2. Same-level cubic interpolation
We develop a new same-level cubic interpolation in HPEZ, which can further improve prediction accuracy. In the traditional interpolation design (Liu et al., 2022; Zhao et al., 2021), at each interpolation level, the neighbor points of each data point to be interpolated are limited on the higher levels (interpolation levels with larger strides). For the 1D cubic spline interpolation applied on a data point with stride $s$, 4 neighbor points with distance $s$ and $3s$ are used, which have been predicted on the higher interpolation levels. As shown in Figure 6 (note that $s$ is the distance between each closest hollow and solid point), the first row shows this interpolation method, in which all the hollow data points (on the current interpolation level) are predicted by the solid data points (on higher interpolation levels). If we are able to include more neighbors for each point (for example, the 2 white points with a distance of $2s$ to it), the prediction accuracy can be improved. As illustrated in the 2nd and 3rd rows of Figure 6, instead of traversing through all the white data points in one step, HPEZ splits the 1D cubic spline interpolation into 2 steps. In the first step (the second row of Figure 6), half of the white points are interpolated by inter-level interpolation (the existing interpolation) with 4 neighbor points. In the second round, the rest half of the white points are interpolated by the same-level interpolation with 6 neighbor points for each, including points interpolated on higher interpolation levels and the current interpolation level. With this new interpolation, half of the data points are predicted with two more neighbor points to achieve better prediction accuracy. Similar to the deductions in Section 5.2, for a data point $p_{i}$, with its 6 neighbor points $d_{i-3}$, $d_{i-2}$, $d_{i-1}$, $d_{i+1}$, $d_{i+2}$, and $d_{i+3}$ the same-level cubic spline interpolation formula would be the following two. Eq. 13 is for the not-a-knot cubic spline and Eq. 14 is for the natural cubic spline. The same strategy can also be extended to the multi-dimensional interpolation, splitting it into 2 steps each with halved data points.
@@ 6.2. Global Interpolation Tuning
  • New in HPEZ: Applying inner-level interpolation or not (only for cubic interpolation);
Similar to QoZ, the sampled data are used for performing compression tests with all the available interpolation configurations. Then, HPEZ selects the interpolation configuration with the lowest average absolute prediction error as the final tuning result.
@@ 6.3. Dynamic Dimension Freezing
The dynamic dimension freezing in HPEZ is designed to avoid inaccurate interpolation predictions along non-smooth dimensions. For a multi-dimensional input data array, it may present fine smoothness along some of its dimensions but present bad smoothness along the other dimensions. In those cases, both the 1D-style and multi-dimensional interpolation will fail in achieving high prediction accuracy as they will involve interpolations along non-smooth directions. The dimension freezing is that, given one dimension, HPEZ sets anchor points along those dimensions with stride 1 (without intervals) and never performs interpolations along those dimensions. Figure 8 uses the interpolation on a 3D data block as an example of dimension freezing. For a clear view, only the 1D interpolations are shown. Figure 8 (a) is the normal 1D interpolations without a frozen dimension, and Figure 8 (b) is the 1D interpolations with a dimension frozen, in which no interpolations are made along the frozen dimension. With this dynamic strategy, HPEZ does not require data smoothness along all dimensions to optimize its compression ratio. According to our experimental results, compared to the highly improved prediction accuracy and greatly reduced quantization bin size, the storage overhead for additional anchor points is affordable. To determine whether to freeze a dimension and which dimension should be frozen, the auto-tuning module of HPEZ first specifies the most non-smooth dimension in the input data array in the statistical analysis (Section 6.1), then separately tunes 2 optimized interpolation configurations with/without this dimension frozen. If freezing this dimension presents a better compression ratio, HPEZ will freeze this dimension.